f(x) = max (x, x2)
Find the points where f(x) is non differentiable.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

f(x) = max(x, x^2)
f(x) = x^2, x < 0
= x, 0 \leq x < 1
= x^2, x\geq 1
So there are two points around which slope is changing.
x = 0, 1
are the points where f(x) is non-differentiable.